# Adaptive Data-Borrowing for Improving Treatment Effect Estimation using External Controls

### v1

@@ Abstract
Randomized controlled trials (RCTs) often exhibit limited inferential efficiency in estimating treatment effects due to small sample sizes. In recent years, the combination of external controls has gained increasing attention as a means of improving the efficiency of RCTs. However, external controls are not always comparable to RCTs, and direct borrowing without careful evaluation can introduce substantial bias and reduce the efficiency of treatment effect estimation. In this paper, we propose a novel influence-based adaptive sample borrowing approach that effectively quantifies the "comparability” of each sample in the external controls using influence function theory. Given a selected set of borrowed external controls, we further derive a semiparametric efficient estimator under an exchangeability assumption. Recognizing that the exchangeability assumption may not hold for all possible borrowing sets, we conduct a detailed analysis of the asymptotic bias and variance of the proposed estimator under violations of exchangeability. Building on this bias-variance trade-off, we further develop a data-driven approach to select the optimal subset of external controls for borrowing. Extensive simulations and real-world applications demonstrate that the proposed approach significantly enhances treatment effect estimation efficiency in RCTs, outperforming existing approaches.
@@ 1 Introduction
Randomized controlled trials (RCTs) are regarded as the gold standard for estimating treatment effects, as randomization effectively eliminates confounding bias imbens2015causal ; Hernan-Robins2020 . However, their inferential efficiency is often limited by small sample sizes due to high costs and lengthy recruitment periods athey2017state . To address this challenge, there has been growing interest in novel clinical trial designs that leverage external real-world datasets containing only control arms, referred to as external controls, to improve treatment effect estimation in RCTs. For example, in digital marketing, platforms may combine small-scale A/B test data with historical single-arm user behavior logs to enhance inference bojinov2023design . In medical research, oncology trials often incorporate historical control data to strengthen causal conclusions hobbs2011hierarchical . Similarly, rare disease studies frequently augment small randomized trials with matched external controls from disease registries to improve statistical power wang2019using .
To fully exploit external controls, most studies rely on the exchangeability assumption, which posits that the distribution of potential outcome under control remains invariant between RCTs and external controls when conditioned on baseline covariates dahabreh2019generalizing ; dahabreh2020extending ; li2023improving . However, the exchangeability assumption requires that individuals in both RCTs and external controls follow the same pattern between the covariates and the potential outcome under control, which may be less plausible in real-world applications due to individual heterogeneity gao2024improving . In practice, external controls usually have significantly larger sample sizes than RCTs and often exhibit greater individual heterogeneity li2023improving ; fda2023considerations . When some individuals in the external controls display patterns that differ from those in the RCTs, the exchangeability assumption is violated and leads to biased conclusions.
@@ 2 Related Work
There has been an increasing amount of research in settings where RCT data are augmented with external controls to improve efficiency in estimating treatment effects since pocock1976combination , termed external control, historical control, or history borrowing in related literature. Under the exchangeability assumption (or its analog) for the potential outcome under control, several approaches dahabreh2019generalizing ; li2023improving ; valancius2024causal have been proposed to estimate treatment effects in RCTs using external controls. However, the exchangeability assumption may be less plausible in practice due to individual heterogeneity. As a result, combining the RCT data and external controls directly will lead to bias. To address the issue, various methods have been developed, including matching and bias adjustment  stuart2008matching , power priors neuenschwander2009note , meta-analytic schoenfeld2019design . More recently, a state-of-the-art approach proposed by gao2024improving revealed the limitations of previous approaches and developed an adaptive lasso-based borrowing approach, which borrows a comparable subset of external controls through bias penalization. Different from these studies, we propose a novel influence-based sample borrowing approach that utilizes influence functions to quantify the perturbation effect of each external control sample on the outcome model in the RCT controls. This enables the derivation of individual-level influence scores that capture the compatibility of external controls. Also, we develop approaches for borrowing the optimal subset of external controls. In this sense, our approach complements existing methodologies.
Beyond combining RCT data with external control data, there are many other settings for data combination (Degtiar-Rose2023, ; kallus2024role, ; Colnet-etal2024, ; Wu-etal-2024-Compare, ). For example, one can combine RCT (or experimental) data with external data that contain only covariates Dahabreh-etal2019 ; Dahabreh-etal2020 . In such settings, the goal is typically to generalize the causal effect from the RCT to the external population, rather than to improve causal effect estimation within the RCT itself. Another common setting involves combining RCT data with confounded observational data that suffer from unobserved confounders Chen-Cai-2021 ; Yang-etal2023 . In addition, several studies have combined multiple datasets for causal inference Han-etal2023-NIPS ; Han-etal-2023 . In addition, several studies have combined multiple datasets for causal inference (Han-etal2023-NIPS, ; Han-etal-2023, ). When combining multiple datasets, it is important not only to consider which identifiability assumptions to adopt in order to improve causal effect estimation, but also to address how to preserve privacy. Unlike these studies that either use the entire external dataset or do not use external data at all, our method performs individual-level selection of external data.
@@ 3 Preliminaries
where $\hat{\mu}_{0}(x)$, $\hat{\mu}_{1}(x)$, and $\hat{e}_{1}(x)$ are estimates of $\mu_{0}(x)$, $\mu_{1}(x)$, and $e_{1}(x)$, respectively. This is the classical augmented inverse probability weighting (AIPW) estimator, widely studied in prior work (Bang-Robins-2005, ; Kang-Schafer-2007, ; Tan2007, ). The estimator $\hat{\tau}_{\textup{aipw}}$ serves as an asymptotically unbiased benchmark for $\tau$.
Nevertheless, estimating $\tau$ using only the RCT data often lacks efficiency due to small sample sizes, resulting from high costs, long recruitment periods, and ethical or feasibility constraints. This paper aims to improve the estimation efficiency of $\tau$ by fully leveraging external controls.
@@ 5 Adaptive Sample Borrowing via Influence Function
Denote $\mathcal{C}=\{X_{i},A_{i}=0,Y_{i},R_{i}=1,i\in\mathcal{E}\}$ as the RCT’s controls, and $N_{\mathcal{C}}$ is the sample size. Given RCT’s controls $\{Z_{i}=(X_{i},Y_{i}),Z_{i}\in\mathcal{C}\}$. To train the model $\hat{\mu}_{0}$ (parameterized by $\theta$), we learn the model parameters by the empirical risk minimizer, $\hat{\theta}\overset{\text{def}}{=}\arg\min_{\theta\in\Theta}\sum_{Z_{i}\in\mathcal{C}}L(Z_{i};\theta),+\frac{\lambda}{2}\|\theta\|_{2}^{2},$ where $L(Z_{i};\theta)=(Y_{i}-\hat{\mu}_{0}(X_{i};\theta))^{2}$ that is twice-differentiable and convex in $\theta$. and $\lambda>0$ controls regularization strength. For an added external control sample $z=(x,y)$, we denote the modified parameters as $\hat{\theta}_{+z}$, which is obtained by retraining the model after adding $z$ as $\hat{\theta}_{+z}\overset{\textup{def}}{=}\arg\min_{\theta\in\Theta}\sum_{Z_{i}\in\mathcal{C}\cup z}L(Z_{i};\theta),+\frac{\lambda}{2}\|\theta\|_{2}^{2}$. Then, we can obtain the influence score of the external control sample $z$ on loss over RCT’s controls,
@@
Unlike the adaptive lasso-based approach, the influence-based approach does not rely on modeling $\mu_{0,\mathcal{O}}(x)$. In addition, it is noteworthy that the influence score is defined at the individual level, and the score for each point is unaffected by other points in the external controls. As a result, it is robust to outliers in external controls. As shown in Figure 1(b), the proposed method effectively selects the top-$K$ points (in red) that are close to the RCT controls, while remaining robust to outliers in the external controls.
@@ 6.2 Find the Optimal Borrowed Set
where $\mathcal{\bf S}=\{\mathcal{S}_{k}:k=1,...,N_{\mathcal{O}}\}$, $\mathcal{S}_{k}$ denotes a subset of external controls corresponding to the $k$ smallest influence scores, $N_{\mathcal{O}}$ is the sample size of external controls.
Importantly, the proposed selection strategy for $\mathcal{S}^{*}$ does not rely on Assumption 2 or 3, nor does it impose any restrictions on the distribution of the external controls. Therefore, our method exhibits high applicability across a variety of settings.
RCT data typically have relatively small sample sizes, as their collection is often time-consuming and costly, and further limited by the inclusion criteria set by the experimenter Kallus-Puli2018 . In comparison, external controls are observational, more readily available, and often drawn from larger and more diverse sources, exhibiting greater individual heterogeneity li2023improving ; fda2023considerations ; colnet2024causal . As a result, it is inevitable that some individuals (or outliers) in the external controls will exhibit patterns that differ significantly from those of the RCT controls. In this case, the adaptive lasso-based approach can only achieve sub-optimal performance (as discussed in Section 4), whereas our proposed method can effectively accommodate such scenarios.
@@ 7 A Synthetic Example
where $X_{\textup{rt}},X_{\textup{ec}}\sim U(0,2)$, and $\epsilon_{\textup{rt}}\sim N(0,0.2)$, $\epsilon_{\textup{ec}}\sim N(0,0.5)$. The $X_{\textup{rt}},X_{\textup{ec}}$ is the 1-dimensional covariates. And the external controls contain five outlier samples.
@@ 8.1 Experimental Setup
As outlined in Table 2, for the “Linear” mechanism, we simulate the RCTs baseline covariates $X_{\text{rt}}\sim N(\mu_{1},\sigma_{1})$ and the external controls baseline covariates $X_{\text{ec}}\sim N(\mu_{2},\sigma_{2})$. Here, $\mu_{1}\neq\mu_{2}$, with $|\mu_{2}-\mu_{1}|$ quantifying the magnitude of covariate shift between the two groups. The standard deviation $\sigma_{2}>\sigma_{1}$ reflects greater heterogeneity in the external controls compared to the RCTs. We initially set $\mu_{1}=0,\sigma_{1}=1$ and $\mu_{2}=0.1,\sigma_{2}=2$. The coefficient $\Delta\beta_{1}\sim U([0.8,1.2]^{d})$ characterizes the structural differences in outcome models between the external controls and the RCTs. $\delta$ and $T$ constitute the effect of concurrency bias, in which $T_{i},i\in\mathcal{O}$, is simulated by taking values of $(0,1,2)$ with probability $1/3$ and $\delta$ represents the level of inconcurrency, where we set $\delta=0.1$. Similarly, for the “Nonlinear” mechanism, $\tilde{X}_{\text{rt}}\sim N_{[-2,2]}(\mu_{1},\sigma_{1})$ and $\tilde{X}_{\text{ec}}\sim N_{[-4,4]}(\mu_{2},\sigma_{2})$, both follow a truncated normal distribution. The $\Delta\beta_{2}$ characterizes the structural differences and $\tilde{\delta}$ represents the level of inconcurrency. We set $\Delta\beta_{2}\sim U([0.8,1.2]^{d})$ and $\tilde{\delta}=1.0$.
@@
$\beta_{1}\sim U([-1,1]^{d}),\epsilon_{rt}\sim N(0,1.0)$
$\epsilon_{ec}\sim N(0,1.5)$
@@
$\beta_{2}\sim U([-1,1]^{d}),\tilde{\epsilon}_{rt}\sim N(0,1.0)$
$\tilde{\epsilon}_{ec}\sim N(0,2.0)$
@@
Baselines and Evaluation metrics. We compare our proposed influence-based approach $(\hat{\tau}_{\textup{if}})$ with the following baselines in the above datasets. (a) the augmented inverse probability weighting (AIPW) estimator without borrowing ($\hat{\tau}_{\textup{aipw}}$cao2009improving (based only on the RCT data); (b) the AIPW estimator with full borrowing ($\hat{\tau}_{\textup{full}}$li2023improving . (c) the AIPW estimator with adaptive lasso-based borrowing approach ($\hat{\tau}_{\textup{lasso}}$gao2024improving . We report three evaluation metrics, including standard deviation, bias, and mean squared errors (MSE), with detailed definitions provided in Section 6.2.

### v2

@@ Abstract
Randomized controlled trials (RCTs) often exhibit limited inferential efficiency in estimating treatment effects due to small sample sizes. In recent years, the combination of external controls has gained increasing attention as a means of improving the efficiency of RCTs. However, external controls are not always comparable to RCTs, and direct borrowing without careful evaluation can introduce substantial bias and reduce the efficiency of treatment effect estimation. In this paper, we propose a novel influence-based adaptive sample borrowing approach that effectively quantifies the “comparability” of each sample in the external controls using influence function theory. Given a selected set of borrowed external controls, we further derive a semiparametric efficient estimator under an exchangeability assumption. Recognizing that the exchangeability assumption may not hold for all possible borrowing sets, we conduct a detailed analysis of the asymptotic bias and variance of the proposed estimator under violations of exchangeability. Building on this bias-variance trade-off, we further develop a data-driven approach to select the optimal subset of external controls for borrowing. Extensive simulations and real-world applications demonstrate that the proposed approach significantly enhances treatment effect estimation efficiency in RCTs, outperforming existing approaches.
@@ 1 Introduction
Randomized controlled trials (RCTs) are regarded as the gold standard for estimating treatment effects, as randomization effectively eliminates confounding bias imbens2015causal ; Hernan-Robins2020 ; Wu-etal2022 . However, their inferential efficiency is often limited by small sample sizes due to high costs and lengthy recruitment periods athey2017state ; hu2023longterm ; yang2024learning ; Wu-etal-ShortLong . To address this challenge, there has been growing interest in novel clinical trial designs that leverage external real-world datasets containing only control arms, referred to as external controls, to improve treatment effect estimation in RCTs. For example, in digital marketing, platforms may combine small-scale A/B test data with historical single-arm user behavior logs to enhance inference bojinov2023design . In medical research, oncology trials often incorporate historical control data to strengthen causal conclusions hobbs2011hierarchical . Similarly, rare disease studies frequently augment small randomized trials with matched external controls from disease registries to improve statistical power wang2019using .
To fully exploit external controls, most studies rely on the exchangeability assumption, which posits that the distribution of potential outcomes under control remains invariant between RCTs and external controls when conditioned on baseline covariates dahabreh2019generalizing ; dahabreh2020extending ; li2023improving ; wu2024comparative ; Wu-Mao2025 . However, the exchangeability assumption requires that individuals in both RCTs and external controls follow the same pattern between the covariates and the potential outcome under control, which may be less plausible in real-world applications due to individual heterogeneity gao2024improving ; Wu-etal-2024-Harm ; Wu-etal-2025-Safe . In practice, external controls usually have significantly larger sample sizes than RCTs and often exhibit greater individual heterogeneity li2023improving ; wu2024comparative ; fda2023considerations . When some individuals in the external controls display patterns that differ from those in the RCTs, the exchangeability assumption is violated and leads to biased conclusions.
@@ 2 Related Work
There has been an increasing amount of research in settings where RCT data are augmented with external controls to improve efficiency in estimating treatment effects since pocock1976combination , termed external control, historical control, or history borrowing in related literature. Under the exchangeability assumption (or its analog) for the potential outcome under control, several approaches dahabreh2019generalizing ; li2023improving ; valancius2024causal have been proposed to estimate treatment effects in RCTs using external controls. However, the exchangeability assumption may be less plausible in practice due to individual heterogeneity. As a result, combining the RCT data and external controls directly will lead to bias. To address the issue, various methods have been developed, including matching and bias adjustment stuart2008matching , power priors neuenschwander2009note , meta-analytic predictive priors schoenfeld2019design , and conformal prediction Zhu-etal-ICML25 . More recently, a state-of-the-art approach proposed by gao2024improving revealed the limitations of previous approaches and developed an adaptive lasso-based borrowing approach, which borrows a comparable subset of external controls through bias penalization. Unlike these studies, we propose a novel influence-based sample borrowing approach that utilizes influence functions to quantify the perturbation effect of each external control sample on the outcome model in the RCT controls. This enables the derivation of individual-level influence scores that capture the compatibility of external controls. Also, we develop approaches for borrowing the optimal subset of external controls. In this sense, our approach complements existing methodologies.
Beyond combining RCT data with external control data, there are many other settings for data combination (wu2024comparative, ; Wu-Mao2025, ; Degtiar-Rose2023, ; kallus2024role, ; Colnet-etal2024, ). For example, one can combine RCT (or experimental) data with external data that contain only covariates Dahabreh-etal2019 ; Dahabreh-etal2020 . In such settings, the goal is typically to generalize the causal effect from the RCT to the external population, rather than to improve causal effect estimation within the RCT itself. Another common setting involves combining RCT data with confounded observational data that suffer from unobserved confounders Chen-Cai-2021 ; Yang-etal2023 . In addition, several studies have combined multiple datasets for causal inference Wu-Mao2025 ; Han-etal2023-NIPS ; Han-etal-2023 . When combining multiple datasets, it is important not only to consider which identifiability assumptions to adopt in order to improve causal effect estimation, but also to address how to preserve privacy. Unlike these studies that either use the entire external dataset or do not use external data at all, our method performs individual-level selection of external data.
@@ 3 Preliminaries
where $\hat{\mu}_{0}(x)$, $\hat{\mu}_{1}(x)$, and $\hat{e}_{1}(x)$ are estimates of $\mu_{0}(x)$, $\mu_{1}(x)$, and $e_{1}(x)$, respectively. This is the classical augmented inverse probability weighting (AIPW) estimator, widely studied in prior work (Bang-Robins-2005, ; Kang-Schafer-2007, ; Tan2007, ; Wu-Han2024, ; Wu-Tan2024, ). The estimator $\hat{\tau}_{\textup{aipw}}$ serves as an asymptotically unbiased benchmark for $\tau$. Nevertheless, estimating $\tau$ using only the RCT data often lacks efficiency due to small sample sizes, resulting from high costs, long recruitment periods, and ethical or feasibility constraints. This paper aims to improve the estimation efficiency of $\tau$ by fully leveraging external controls.
@@ 5 Adaptive Sample Borrowing via Influence Function
Denote $\mathcal{C}=\{X_{i},A_{i}=0,Y_{i},R_{i}=1,i\in\mathcal{E}\}$ as the RCT’s controls, and $N_{\mathcal{C}}$ is the sample size. Given RCT’s controls $\{Z_{i}=(X_{i},Y_{i}),Z_{i}\in\mathcal{C}\}$. To train the model $\hat{\mu}_{0}$ (parameterized by $\theta$), we learn the model parameters by the empirical risk minimizer, $\hat{\theta}\overset{\text{def}}{=}\arg\min_{\theta\in\Theta}\sum_{Z_{i}\in\mathcal{C}}L(Z_{i};\theta),+\frac{\lambda}{2}\|\theta\|_{2}^{2},$ where $L(Z_{i};\theta)=(Y_{i}-\hat{\mu}_{0}(X_{i};\theta))^{2}$ that is twice-differentiable and convex in $\theta$, and $\lambda>0$ controls regularization strength. For an added external control sample $z=(x,y)$, we denote the modified parameters as $\hat{\theta}_{+z}$, which is obtained by retraining the model after adding $z$ as $\hat{\theta}_{+z}\overset{\textup{def}}{=}\arg\min_{\theta\in\Theta}\sum_{Z_{i}\in\mathcal{C}\cup z}L(Z_{i};\theta),+\frac{\lambda}{2}\|\theta\|_{2}^{2}$. Then, we can obtain the influence score of the external control sample $z$ on loss over RCT’s controls,
@@
Unlike the adaptive lasso-based approach, the influence-based approach does not rely on modeling $\mu_{0,\mathcal{O}}(x)$. In addition, it is noteworthy that the influence score is defined at the individual level, and the influence score for each point is unaffected by other points in the external controls. As a result, it is robust to outliers in external controls. As shown in Figure 1(b), the proposed method effectively selects the top-$K$ points (in red) that are close to the RCT controls, while remaining robust to outliers in the external controls.
@@ 6.2 Find the Optimal Borrowed Set
where $\mathcal{\bf S}=\{\mathcal{S}_{k}:k=1,...,N_{\mathcal{O}}\}$, $\mathcal{S}_{k}$ denotes a subset of external controls corresponding to the top-$k$ smallest influence scores, $N_{\mathcal{O}}$ is the sample size of external controls.
Importantly, the proposed selection strategy for $\mathcal{S}^{*}$ does not rely on Assumption 2 or 3, nor does it impose any restrictions on the distribution of the external controls. Therefore, our method exhibits high applicability across a variety of settings. RCT data typically have relatively small sample sizes, as their collection is often time-consuming and costly, and further limited by the inclusion criteria set by the experimenter Kallus-Puli2018 . In comparison, external controls are observational, more readily available, and often drawn from larger and more diverse sources, exhibiting greater individual heterogeneity li2023improving ; fda2023considerations ; colnet2024causal . As a result, it is inevitable that some individuals (or outliers) in the external controls will exhibit patterns that differ significantly from those of the RCT controls. In this case, the adaptive lasso-based approach can only achieve sub-optimal performance (as discussed in Section 4), whereas our proposed method can effectively accommodate such scenarios.
@@ 7 A Synthetic Example
where $X_{\textup{rt}},X_{\textup{ec}}\sim U(0,2)$, and $\epsilon_{\textup{rt}}\sim N(0,0.2^{2})$, $\epsilon_{\textup{ec}}\sim N(0,0.5^{2})$. The $X_{\textup{rt}},X_{\textup{ec}}$ is the 1-dimensional covariates. And the external controls contain five outlier samples.
@@ 8.1 Experimental Setup
As outlined in Table 2, for the “Linear” mechanism, we simulate the RCTs baseline covariates $X_{\text{rt}}\sim N(\mu_{1},\sigma_{1}^{2})$ and the external controls baseline covariates $X_{\text{ec}}\sim N(\mu_{2},\sigma_{2}^{2})$. Here, $\mu_{1}\neq\mu_{2}$, with $|\mu_{2}-\mu_{1}|$ quantifying the magnitude of covariate shift between the two groups. The standard deviation $\sigma_{2}>\sigma_{1}$ reflects greater heterogeneity in the external controls compared to the RCTs. We initially set $\mu_{1}=0,\sigma_{1}=1$ and $\mu_{2}=0.1,\sigma_{2}=2$. The coefficient $\Delta\beta_{1}\sim U([0.8,1.2]^{d})$ characterizes the structural differences in outcome models between the external controls and the RCTs. $\delta$ and $T$ constitute the effect of concurrency bias, in which $T_{i},i\in\mathcal{O}$, is simulated by taking values of $(0,1,2)$ with probability $1/3$ and $\delta$ represents the level of inconcurrency, where we set $\delta=0.1$. Similarly, for the “Nonlinear” mechanism, $\tilde{X}_{\text{rt}}\sim N_{[-2,2]}(\mu_{1},\sigma_{1}^{2})$ and $\tilde{X}_{\text{ec}}\sim N_{[-4,4]}(\mu_{2},\sigma_{2}^{2})$, both follow a truncated normal distribution. The $\Delta\beta_{2}$ characterizes the structural differences and $\tilde{\delta}$ represents the level of inconcurrency. We set $\Delta\beta_{2}\sim U([0.8,1.2]^{d})$ and $\tilde{\delta}=1.0$.
@@
$\beta_{1}\sim U([-1,1]^{d}),\epsilon_{rt}\sim N(0,1.0^{2})$
$\epsilon_{ec}\sim N(0,1.5^{2})$
@@
$\beta_{2}\sim U([-1,1]^{d}),\tilde{\epsilon}_{rt}\sim N(0,1.0^{2})$
$\tilde{\epsilon}_{ec}\sim N(0,2.0^{2})$
@@
Baselines and Evaluation metrics. We compare our proposed influence-based approach $(\hat{\tau}_{\textup{if}})$ with the following baselines in the above datasets. (a) the augmented inverse probability weighting (AIPW) estimator without borrowing ($\hat{\tau}_{\textup{aipw}}$) cao2009improving (based only on the RCT data); (b) the AIPW estimator with full borrowing ($\hat{\tau}_{\textup{full}}$) li2023improving . (c) the AIPW estimator with adaptive lasso-based borrowing approach ($\hat{\tau}_{\textup{lasso}}$) gao2024improving . We report three evaluation metrics, including standard deviation (std), bias, and mean squared errors (MSE), with detailed definitions provided in Section 6.2.